# LetsGo: Large-Scale Garage Modeling and Rendering via LiDAR-Assisted Gaussian Primitives

### v1

@@ 3.1 Raw Data Acquisition
Scanning and modeling a large garage is a non-trivial task. Underground and indoor garages often face challenges in receiving GPS signals due to the physical barriers presented by the structures and materials surrounding them, making camera pose estimation for scanning and modeling difficult. Furthermore, there are always large-scale texture-less regions inside a garage, e.g., floors and walls. The parked vehicles often contain transparent glasses, and their surfaces are sometimes reflective. This complicated feature matching between images for camera pose estimation and 3D geometry reconstruction. Using a LiDAR sensor for scanning and modeling can provide detailed geometric information. However, the RGB color for each scanned point is not associated.
@@
Capturing Device. To address these problems, we design a lightweight handheld scanning device named “Polar” to jointly collect color and geometric information about the garage. The data collection unit of the Polar device comprises a color fisheye camera, a LiDAR sensor, and an IMU sensor, as visualized in Fig. 2. The fisheye camera captures RGB color information. It has a resolution of 6K and a field of view (FOV) of $270\times 360$ degrees, allowing for quick and comprehensive recording of vivid color data. The LiDAR sensor collects 3D point clouds, recording geometric information at a rate of 2.6 million points per second. With a measurement accuracy of 1 to 1.5 cm and a maximum detection distance of 50 m, it is ideal for 3D scanning of large garage scenes. The IMU sensor provides acceleration information on the device’s motion, enabling more accurate pose estimation. In addition to the data acquisition unit, we also equip the Polar device with a data processing unit consisting of a mini PC for real-time SLAM calculations. The data processing unit is powered by two removable batteries that provide more than 30 minutes of single scan endurance. With this unit, one can use a smartphone to connect with the Polar device and preview 3D point cloud reconstruction results in real-time. The entire Polar device weighs no more than 1kg, making it suitable for handheld scanning applications.
Table 1: Detailed illustrations of our GarageWorld dataset.
Campus 1 Underground Flat & Sloped Surfaces 38447.86 8479 1.9B 121.9M Great
Campus 2 Underground Flat & Sloped Surfaces 28046.37 7772 1.4B 95.8M Great
Shopping Mall 1 Indoor (Multi-level) Spiral & Circular Paths 32646.68 5792 1.14B 40.78M Fair
Shopping Mall 2 Outdoor Spiral & Sloped paths 13495.92 2280 0.6B 77.1M Good
Office Building Indoor With Mechanical Parking System 22159.25 9308 1.15B 72.3M Good
@@
Scanning Scheme. We use the Polar device to scan various garages for modeling and free-viewpoint rendering. For each garage, our scanning trajectory aligns with the common vehicle trajectory when driving through garages. We collect data along each trajectory four times: one for forward-facing data collection, one for backward-facing, one for side-left, and one for side-right, resulting in comprehensive capturing of the garage. We set the camera to auto exposure and auto ISO to accommodate complex lighting changes in the garage. For garages that are partially open-air and partially covered, the data is collected when the sunlight is weak, i.e., during the early morning or evening. This ensures the images captured at the transition between open-air and covered areas share similar illumination. We also apply a short pause for data capturing at the transition area, allowing the sensor to adapt to different lighting conditions and thus ensuring the images are neither overexposed nor underexposed. For garages with motion-activated lights, we ensure the data is collected after the light turns on, maintaining consistent lighting conditions across all images. The travel speed for data collection is at around 1.0±0.2 m/s, with a turning speed of 15°±3 °/s. In areas with insufficient light, these parameters are reduced to 0.5 m/s and 10 °/s, respectively, to avoid motion blur. Some garages are very large, for example, over 30,000 square meters. Thus, we divide the large garage into small subsections and collect data for each subsection. The data is fused later after collection.
@@ 3.2 GarageWorld Dataset
We collect data for five garages, including two underground garages on Campus, one indoor garage with multi-levels at Shopping Mall One, one outdoor surface parking at Shopping Mall Two, and one indoor garage with a mechanical parking system in an Office Building. These garages comprise various challenging structures, such as sloped surfaces with distorted lines, internal circular or spiral paths, vehicle elevators, etc., as shown in Fig.  11. Almost all our scanned garages contain electric vehicle charging stations, making them eco-friendly. Tab. 1 provides an overall description of the garages.
Underground Garages. We first collect data for two underground garages on campus. This design for constructing garages beneath the ground is often employed in areas with limited above-ground space, providing a solution that preserves surface area for other uses. The coverage of the two garages is around 38.4k and 28k square meters, respectively, and they only have one floor for vehicle parking with flat and sloped surfaces. The two garages are illuminated by constant light conditions, with regular fluorescent light tubes on the ceiling. We collect 8,479 and 7,772 images for the two garages, respectively, with 1.9 and 1.4 billion point clouds.
@@
Indoor Garage with Multi-floors. Staking garages to multi-floors is a common design in dense urban environments. We collect data for this type of parking garage in Shopping Mall One. The different floors of this garage are connected by spiral and circular paths, which have a semi-open structure and are partially illuminated by Sunlight and partially by indoor lights. Given that the lights in the shopping mall are not turned on in the early morning, we collect data for this garage during the early evening to maintain consistent illumination between different images. The total coverage for this garage data is over 32k square meters, with 5,792 images and a point cloud containing 1.14 billion points.
@@
Outdoor Parking. The outdoor surface parking facility is often in areas with large spaces or on the top of a commercial building. We collect data for a garage with this type located on the top of Shopping Mall Two. The entrance to this parking space is from indoor to outdoor, containing spiral and sloped paths. During daylight time, the outdoor illumination is stronger than indoor. At night, the limited dim streetlights are insufficient for photography requirements. Therefore, we conduct our data capture during the early morning and evening when the sunlight is soft, and the illumination between indoors and outdoors is similar. We scanned an area surpassing 13,000 square meters for this garage, resulting in 2,280 images and a point cloud containing 0.6 billion points.
Indoor Garage with a Mechanical Parking System. We also scan several indoor garages with a mechanical parking system in an office building. These garages are typically small in size and contain colored surfaces. The total coverage of these garages is around 22k square meters, containing 9,308 images and 1.15 billion points in point clouds.
3.3 Relative Pose Estimation and Mesh Reconstruction
We start processing the collected garage data by Polar device calibration and dynamic object removal, followed by relative pose estimation for the images captured at different time steps and mesh reconstruction of the entire garage. The detailed descriptions for each step are presented below.
For the intrinsic parameter calibration of the fisheye camera and the IMU sensor, we leverage the program provided by OpenCV and the Allan Variance ROS toolbox [51, 17, 16, 39, 45], respectively. The extrinsic calibration between the fisheye camera and the IMU is done using the Kalibr calibration program [51]. Due to the sparsity, noisiness, and uncolored nature of the point cloud data collected by our LiDAR sensor, it is hard to establish correspondences between the LiDAR data and the images captured by the fisheye camera, making relative pose estimation between the sensors difficult. To address this issue, we introduce an additional sensor, a FARO laser scanner, which provides dense, accurate, and colored point clouds and thus functions as a bridge for the relative pose estimation between the LiDAR sensor and the fisheye camera. The detailed calibration steps between the sensors are presented in the supplementary material.
Figure 3: Overview of our LiDAR-assisted Gaussian splatting framework. Initially, we generate a base mesh using color and depth data collected by our self-designed Polar device. Then, the data is partitioned into blocks for parallel and rapid processing. Next, we utilize high-quality scanned point clouds as the input for our LiDAR-GS (Sec.4.1) algorithm. Besides photometric supervision, we also apply our novel unbiased Gaussian depth regularizer for geometric supervision. Finally, our system produces photorealistic rendering results.
To remove the dynamic objects in the collected data, we apply Segment Anything [27] to the images. For the point cloud data, we estimate whether the points scanned at one LiDAR frame are also observed at other frames, similar to Schauer and Nüchter [54].
After this preprocessing step, we employ the LiDAR-Inertial-Visual(LIV) SLAM [59] to estimate the relative poses of the sensor between different time steps. The LIV-SLAM system integrates a tightly coupled LiDAR-Inertial Odometry and Visual-Inertial Odometry, along with a joint optimization approach between LiDAR and camera data, for relative pose estimation. Specifically, we employ LIO to fuse IMU data with point clouds collected at different time steps by the LiDAR sensor. This process yields accurate distance measurements and effectively mitigates drift in the IMU data. Subsequently, using the IMU data updated by LIO, we implement VIO among captured fisheye images to estimate their relative poses. Following this, we utilize joint factor graph optimization [35] to refine the estimated poses of the LiDAR and camera. The initial pose graph is constructed by incorporating the updated LiDAR, IMU, and camera poses. Then, bundle adjustment is applied to further optimize the graph. This pose estimation system capitalizes on the strengths of each sensor: IMU data offers short-term motion estimation, LiDAR provides accurate distance measurements, and visual sensors enhance pose estimation in feature-rich environments. Thus, it ensures robust and accurate pose estimation. After merging the point cloud data collected at different time steps using the estimated relative pose, we apply Poisson Reconstruction [24] to convert the point cloud data into an uncolored mesh. To ensure the accuracy and integrity of the resulting mesh, we also compare the reconstructed meth to the original point cloud data to remove incorrect faces.
@@ 3.3 Relative Pose Estimation and Mesh Reconstruction
This geometry-based mesh reconstruction method performs well for Lambertian surface reconstruction. However, it faces challenges for reflective and transparent surfaces, e.g., vehicle glass windows, which is especially common in garage environments. Furthermore, the granularity and preciseness of the reconstructed meshes depend on the density of the 3D points.
4 LiDAR-Assisted Gaussian Primitives
@@
Our Polar scanner with calibrated IMU, LiDAR, and fisheye camera sensors addresses this challenge. The strengths of the different sensors are combined to achieve accurate camera poses and 3D point cloud representations. In the following sections, we describe technical details of our LiDAR-assisted Gaussian splatting with a depth regularizer and a hybrid representation that combines the merits of mesh and 3D Gaussians, followed by training details and experimental evaluation. Fig.3 provides an overview of our LiDAR-assisted Gaussian splatting framework.
4.1 LiDAR-Assisted Gaussian Splatting
@@
Figure 4: Our LiDAR-RGS system comprises a base mesh, a series of residual Gaussians, and a blending weight map. These components enable our system to maintain a photorealistic appearance comparable to LiDAR-GS while significantly reducing runtime memory. This approach strikes an optimal balance between memory consumption and visual fidelity.
@@
In this paper, we use our Polar device with a LiDAR sensor to scan point clouds of the garages. Considering the originally scanned point clouds contain noises, we resample a set of new points from the reconstructed mesh with a uniform sampling strategy. These resampled points, in conjunction with the camera parameters, are used to train the 3DGS representations.
In addition to the original Gaussian splatting, we further introduce a depth-regularizer for the 3DGS training, which leverages the high-fidelity LiDAR data to incorporate depth priors during training. We denote this method as LiDAR-GS, and the original 3DGS method with LiDAR-assisted point cloud for Gaussian initialization as 3DGS.
Depth Regularizer.
Inspired by the depth calculation from NeRF [40], we utilize the rasterization pipeline of Gaussians to compute the depth of each Gaussian primitive:
@@
where $D_{\text{G}}$ is the rendered depth of the 3D Gaussian primitive and $d_{i}$ is the depth of each Gaussian splat in camera perspective.
@@
where $\mathbf{p}=[p_{0},p_{1},p_{2}]$ represents the position of the Gaussian center in the ray space, and $\mathbf{\Sigma}$ is the $3\times 3$ covariance matrix. $\mathbf{(\cdot)}_{m,n}$ represents the corresponding element in the matrix. For detailed derivation and understanding of this process, please refer to the supplementary materials provided with the paper.

### v2

@@ 3.1 Raw Data Acquisition
Scanning and modeling a large garage is a non-trivial task. Underground and indoor garages often face challenges in receiving GPS signals due to the physical barriers presented by the structures and materials surrounding them, making camera pose estimation for scanning and modeling difficult. Furthermore, there are always large-scale texture-less regions inside a garage, e.g., floors and walls. The parked vehicles often contain transparent glasses, and their surfaces are sometimes reflective. This complicates feature matching between images for camera pose estimation and 3D geometry reconstruction. Using a LiDAR sensor for scanning and modeling can provide detailed geometric information. However, the RGB color for each scanned point is not associated.
@@
Capturing Device. To address these problems, we design a lightweight handheld scanning device named “Polar” to jointly collect color and geometric information of the garage. The data collection unit of the Polar device comprises a color fisheye camera, a LiDAR sensor, and an IMU sensor, as visualized in Fig. 2. The fisheye camera captures RGB color information in 30 FPS. It has a resolution of 6K and a field of view (FOV) of $180\times 180$ degrees, allowing for quick and comprehensive recording of vivid color data. The LiDAR sensor collects 3D point clouds, recording geometric information at a rate of 2.6 million points per second. With a measurement accuracy of 1 to 1.5 cm and a maximum detection distance of 50 m, it is ideal for 3D scanning of large garage scenes. The IMU sensor provides acceleration information on the device’s motion, enabling more accurate pose estimation. In addition to the data acquisition unit, we also equip the Polar device with a data processing unit consisting of a mini PC for real-time SLAM calculations. The data processing unit is powered by two removable batteries that provide more than 30 minutes of single scan endurance. With this unit, one can use a smartphone to connect with the Polar device and preview 3D point cloud reconstruction results in real-time. More information about Polar device along with the calibration process is provided in Appendix.A.
@@
Scanning Scheme. We use the Polar device to scan various garages for modeling and free-viewpoint rendering. The scanning trajectory for each garage is meticulously designed to emulate drive-through and parking. We collect data along each trajectory four times: one for forward-facing, one for backward-facing, one for side-left, and one for side-right, ensuring comprehensive capturing of the garage. We set the camera to auto exposure and auto ISO to accommodate complex lighting changes in the garage. For garages that are partially open-air and partially covered, the data is collected when the sunlight is weak, i.e., during the early morning or evening. This ensures the images captured in the transition between open-air and covered areas share similar illumination. We also apply a short pause for data capturing at the transition area, allowing the sensor to adapt to different lighting conditions and thus ensuring the images are neither overexposed nor underexposed. For garages with motion-activated lights, we ensure the data is collected after the light turns on, maintaining consistent lighting conditions across all images. The travel speed for data collection is at around 1.0±0.2 m/s, with a turning speed of 15°±3 °/s. To avoid motion blur in areas with insufficient light, these parameters are reduced to 0.5 m/s and 10 °/s, respectively. Some garages are very large, for example, over 30,000 square meters. Thus, we divide the large garage into small subsections and collect data for each subsection. The data is fused later after collection.
Figure 3: Overview of our LiDAR-assisted Gaussian splatting framework. Initially, we generate a base mesh using color and depth data collected by our self-designed Polar device. The data is then partitioned into blocks for parallel and rapid processing. Next, we downsample the high-quality scanned point clouds into multi-resolution point cloud for our LOD-LiDAR-RGS (Sec.4.2) method initialization. In addition to photometric supervision, we apply our novel unbiased Gaussian depth regularizer (Sec.4.1) for geometric supervision. Finally, our system produces photorealistic LOD rendering results based on the optimized multi-resolution Gaussian representation.
@@ 3.2 GarageWorld Dataset
We collect data for eight garagesAll garage data is captured with the necessary permits., including six underground garages with various types of inside geometries, one indoor garage with multi-floors at Shopping Mall One, and one outdoor surface parking at Shopping Mall Two. We plan to share the datasets with the community for further research. These garages comprise various challenging structures, such as sloped surfaces with distorted lines, internal circular or spiral paths, vehicle elevators, etc., as shown in Fig. 5. Tab. 1 provides an overall description of the garages.
Underground Garages. Most of our captured garages are located underground. This architectural choice is prevalent in regions where above-ground space is scarce, allowing the conservation of surface area for alternative applications. Our underground garages on campus feature a single parking level, characterized by flat and sloped surfaces, and are lit by regular fluorescent light tubes affixed to the ceilings. In contrast, the garages at Shopping Mall Three and the Arts Center have vibrant design elements and superior lighting conditions. Additionally, we conduct a scan of a compact indoor garage within an office building, equipped with a mechanical parking system and featuring colored surfaces.
@@
Indoor Garage with Multi-floor. Staking garages to multi-floor is a common design in dense urban environments. We collect data for this type of parking garage in Shopping Mall One. The different floors of this garage are connected by spiral and circular paths, which have a semi-open structure and are partially illuminated by sunlight and partially by indoor lights. Given that the lights in the shopping mall are not turned on in the early morning, we collect data for this garage during the early evening to maintain consistent illumination between different images.
@@
Outdoor Parking. The outdoor surface parking facility is often in areas with large spaces or on the top of a commercial building. We collect data for a garage with this type located on the top of Shopping Mall Two. The entrance to this parking space is from indoor to outdoor, containing spiral and sloped paths. During daylight time, the outdoor illumination is stronger than indoor. At night, the limited dim streetlights are insufficient for photography requirements. Therefore, we conduct our data capture during the early morning and evening when the sunlight is soft, and the illumination between indoors and outdoors is similar.
3.3 Initial Mesh Reconstruction
We employ the off-the-shell LiDAR-Inertial-Visual(LIV) SLAM [69] to estimate the relative poses of the sensor between different time steps. The LIV-SLAM system integrates a tightly coupled LiDAR-Inertial Odometry and Visual-Inertial Odometry, along with a joint optimization approach between LiDAR and camera data, for relative pose estimation. The pose estimation system leverages the unique capabilities of each sensor of our Polar device to enhance overall accuracy and robustness. The IMU delivers reliable short-term motion estimates, while the LiDAR contributes precise distance measurements. Visual sensors complement these by enriching pose estimation in environments abundant with visual features, ensuring a robust and precise outcome.
@@ 3.3 Initial Mesh Reconstruction
After merging the point cloud data collected at different time steps using the estimated relative pose, we apply Poisson Reconstruction [32] to convert the point cloud data into a mesh. To ensure the accuracy and integrity of the resulting mesh, we also compare the reconstructed meth to the original point cloud data to remove incorrect faces (More details in Appendix.B). This geometry-based mesh reconstruction method performs well for Lambertian surface reconstruction. However, it faces challenges for reflective and transparent surfaces, e.g., vehicle glass windows, which is especially common in garage environments.
4 LiDAR-Assisted LoD Gaussian
@@
Our Polar scanner, equipped with calibrated IMU, LiDAR, and fisheye camera sensors, effectively addresses this challenge. By integrating the unique strengths of these diverse sensors, our system achieves precise camera localization and detailed 3D point cloud generation. In the following sections, we first delve into the technical details of our novel LiDAR-assisted Gaussian splatting technique (Sec. 4.1), which incorporates a custom depth regularizer. By leveraging geometry priors derived from our RGBD scans, this method enhances the realism of rendered images and minimizes the occurrence of floater artifacts. Subsequently, we introduce an innovative framework incorporating Level-of-Detail (LOD) technology into our Gaussian training and rendering process (Sec. 4.2). This framework utilizes a multi-resolution point cloud as its input, significantly boosting rendering speeds. Moreover, we show how it can facilitate rendering expansive garage environments on lightweight devices through a specialized web-based renderer (Sec. 4.3). Fig.3 provides an overview of our LiDAR-assisted Gaussian splatting framework.
4.1 Gaussian Splatting with LiDAR Inputs
@@
Since conventional SfM approaches fail in large garage environments due to hard correspondence establishment between textureless and transparent regions, we use our Polar device with a LiDAR sensor to scan point clouds of the garages. Considering the originally scanned point clouds contain noises, we resample a set of new points from the reconstructed mesh with a uniform sampling strategy. These resampled points, in conjunction with the camera parameters, are used to train the 3DGS representations.
In addition to the image reconstruction loss in the original Gaussian splatting, we further introduce a depth-regularizer for our LiDAR-assisted 3DGS training, which incorporates depth priors derived from the high-quality LiDAR data during training. We denote this method as LiDAR-GS and the original 3DGS method with LiDAR-assisted point cloud for Gaussian initialization as 3DGS.
Depth Regularizer. Inspired by the depth calculation from NeRF [51], we utilize the rasterization pipeline of Gaussians to compute the depth of each Gaussian primitive:
@@
where $D_{\text{G}}$ is the rendered depth and $d_{i}$ is the depth of each Gaussian splat in camera perspective.
@@
where $\mathbf{p}=[p_{0},p_{1},p_{2}]$ represents the position of the Gaussian center in the ray space, and $\mathbf{\Sigma}$ is the $3\times 3$ covariance matrix. $\mathbf{(\cdot)}_{m,n}$ represents the corresponding element in the matrix. For detailed derivation and understanding of this process, please refer to Appendix.C.
@@
Figure 4: Illumination of our LOD-LiDAR-GS within an Octree Structure, accompanied by an LOD rendering strategy. Distance varies from near to far, with Gaussian level resolution transitioning from fine to coarse. The bottom row illustrates the level composition in the rendered view.
4.2 Multi-Resolution Representation
The original 3D Gaussian representation requires a vast number of 3D Gaussians, making it resource-intensive and inefficient for lightweight devices. However, simultaneously loading all Gaussians of the entire scene is inefficient and unnecessary for rendering a specific view.
Inspired by advanced rendering techniques [66] for massive LiDAR point clouds, we introduce a multi-resolution Gaussian framework called LOD-LiDAR-GS. This framework integrates the Level-of-Detail (LOD) rendering to the LiDAR-GS, making it suitable for various devices. We represent the 3D scene with Gaussians at different resolution levels. Each Gaussian has an additional LOD value, alongside its original attributes like 3D position, spherical harmonic coefficients, and opacity. The LOD value dynamically adjusts the cloning and splitting threshold during training. This multi-resolution representation enables fast and lightweight rendering on web-based devices.
Figure 5: Visualization of examples from our GarageWorld dataset.
Multi-Resolution Gaussian Initialization.
We first construct multi-resolution point cloud data for Gaussian initialization. Our Polar device generates dense point clouds with superior accuracy and consistency, even in challenging garage conditions. We use a spacing attribute, $\tau$, to represent the resolution of a point cloud, defined as the minimal distance between points. The finest resolution corresponds to the original point cloud sampled from our reconstructed mesh at the lowest $\tau$. We then increase $\tau$ to $2\tau$ and downsample the point cloud to derive a coarser point cloud, repeating until the point count drops below a threshold, $\epsilon_{p}$. Points in the finest level get an LOD value of $N=L-1$, and those in the coarsest level get an LOD value of $N=0$. We set $\tau$ to 4.0 cm and $\epsilon_{p}$ to 10,000 points. The downsampling operation employs an approximate Poisson-disk sampling method from the PotreeConverter library [66].
LOD-LiDAR-GS Training.
In contrast to LiDAR-GS, our LOD-LiDAR-GS framework processes a multi-resolution point cloud, wherein each resolution level has a distinct LOD value. We devise a new training strategy to handle the complexity caused by multi-resolution input effectively. We initialize a Gaussian model for each resolution level, which operates independently during optimization. Following the original 3DGS [34], the decision to clone or split a Gaussian is based on threshold values for position gradients $\sigma_{pos}$ and variance $\sigma_{var}$. We introduce a scaling factor, $s$, which varies with the LOD value $k$ of each level, modifying the thresholds as $s\cdot\sigma_{pos}$ and $s\cdot\sigma_{var}$:
$s_{k}=\min(\beta_{s}^{L-1-k},s_{max}),\quad k\in[0,L-1]$ (7)
where $\beta_{s}$ and $s_{max}$ are hyperparameters set to $\sqrt{2}$ and $4.0$ respectively.
In our approach, the Gaussian representation at the lower resolution levels is employed to encapsulate the low-frequency content of the scene, while higher resolution levels utilize finer Gaussians to capture the intricate high-frequency details. To use our multi-resolution Gaussian models for image rendering, we adopt a technique analogous to LOD rendering used in traditional computer graphics pipelines. Gaussian subsets from each resolution level are selected based on the rendering camera’s viewing frustum and the projected depth $d$ of each Gaussian. The appropriate Gaussian level $L^{(d)}$ is determined as:
$L^{(d)}=\text{clamp}(\lfloor L^{1-d/d_{max}}\rfloor,0,L-1)$ (8)
where $d_{max}$ is the maximum depth value projected onto the training viewpoint from the point cloud.
During training, we select portions of Gaussians from different resolution levels based on depth ranges to accurately predict the image. We compute the loss by comparing model predictions with ground truth images. Our random-resolution-level (RRL) training strategy renders each training sample with a 50% probability using either the LOD strategy from multiple levels or a single Gaussian level, preventing overfitting and biases in camera pose distribution.
Input : A set of multi-resolution Gaussians stored in .ply files
Output : Gaussian in Octree structure stored in .bin files
// Initialization
Create an octree with depth $L$
Compute bounding box $B$ that covers all Gaussian levels
Initialize $queue=\{\text{root of octree},B,\text{lvl}=0\}$
// First loop: Distribute Gaussians into octree nodes
while $\text{length}(queue)>0$ do
       $currentNode\leftarrow queue.pop()$
       foreach childNode of currentNode do
             Assign bounding box $B_{child}$ and $lvl=\{currentNode.lvl\}+1$ to childNode
             queue.push({childNode, $B_{child}$, $lvl$})
       end foreach
      Read Gaussians from $level\_\{currentNode.lvl\}.ply$
       Distribute Gaussians within $currentNode.B$
end while
Initialize $queue=\{\text{root of octree},B,\text{lvl}=0\}$
$byteOffset,byteSize=0,0$
// Second loop: Prune empty nodes and store data
while $\text{length}(queue)>0$ do
       $currentNode\leftarrow queue.pop()$
       foreach childNode of currentNode do
             if childNode.numPoints ¿ 0 then
                   queue.push({childNode, childNode.B, lvl})
             end if
       end foreach
      $byteSize=\text{size}(currentNode.points)$
       Flush $currentNode.points$ into octree.bin file
       Flush $currentNode.B$, $currentNode.numPoints$, $byteOffset$, $byteSize$ into hierarchy.bin file
       $byteOffset\leftarrow byteOffset+byteSize$
end while
ALGORITHM 1 Octree Construction for LOD-LiDAR-GS
4.3 Web-based Lightweight Renderer
With the proposed multi-resolution Gaussian representation, our LOD-LiDAR-GS significantly enhances rendering efficiency. Instead of using a large number of 3D Gaussians for each image, LOD-LiDAR-GS dynamically selects Gaussian subsets from different resolution levels based on depth ranges. This reduces the complexity of sorting, projection, and accumulation operations, thereby accelerating rendering speed. On GPU systems such as the NVIDIA RTX3090, we implement a 3D Gaussian viewer based on the original 3DGS SIBR viewer. By integrating our LOD rendering strategy, as detailed in Equation 8, we achieve approximately a 4x increase in rendering speed. Additionally, we introduce a lightweight renderer employing an on-demand rendering approach. Unlike mainstream lightweight Gaussian renderers [39, 18], which load all Gaussians into VRAM simultaneously, our renderer dynamically loads only the necessary data into memory according to the current viewpoint and LOD strategy. This method avoids the need for full data loading, enabling the rendering of large-scale environments on web platforms.